OpenStudy (anonymous):

A potential difference of 1400V is established between two parallel plates 4.00cm apart. An electron is released from the negative plate at the same instant that a proton is released from the positive plate. How far from the positive plate will they pass each other ?

OpenStudy (anonymous):

 the electric forces magnitude equal both elecron and proton but, because of mass of the proton is bigger than mass of the electron, they will meet near positive plate. you can find E by using E=V*d you can use F=m*a and then find acceleration both of them then use kinematic equations.